# A Review on Knowledge Graphs for Healthcare: Resources, Applications, and Promises

### v4

@@ 2.3. HKG Utilization
Model-free Utilization. Various query languages can be used for KGs, such as SPARQL, Cypher, and GraphQL (Wang et al., 2020). These query languages allow users to query healthcare KGs using a standardized syntax, thus enabling users to retrieve, manipulate, and analyze data in a structured and consistent way. More complex applications can be further supported by graph queries. For instance, automatic healthcare question answering can be tackled by Natural Language Question-to-Query (NLQ2Query) approach (Kim et al., 2022), where natural language questions are first translated into executable graph queries and then answered by the query responses. HKGs can also be utilized as an up-to-date and trustworthy augmentation to large language models (LLMs) for many applications. Some pioneering studies (Liu et al., 2022c; Guu et al., 2020; Xu et al., 2023; Shi et al., 2023) show that retrieved knowledge triples can improve the reliability of LLMs in various knowledge-intensive tasks, by addressing the nonsensical or unfaithful generation. Moreover, KGs can be a useful tool for fact-checking (Tchechmedjiev et al., 2019; Vedula and Parthasarathy, 2021; Mayank et al., 2022) as they provide a structured representation of information that can be used to quickly and efficiently verify the accuracy of claims. Researchers have explored the utility of HKGs in identifying ingredient substitutions of food (Shirai et al., 2021), COVID-19 fact-checking (Mengoni and Yang, 2022), etc.
@@
Model-based Utilization. Utilizing HKGs in complex reasoning tasks often involves utilizing machine learning models. HKG embeddings (Yu et al., 2021; Su et al., 2022) have shown great potential to tackle these tasks. In particular, HKG embedding models are a class of machine learning models that aim to learn low-dimensional vector representations, or embeddings, of the entities and relations in a knowledge graph. After obtaining HGK embeddings, they can be plugged into any kind of deep neural network and further fine-tuned toward downstream objectives. On the other hand, symbolic logic models represent another prominent approach for KG reasoning due to their interpretability. More specifically, symbolic reasoning models first mine logical rules from existing knowledge by inductive logic programming (Muggleton, 1992), association rule mining (Galárraga et al., 2013), or Markov logic networks (Kok and Domingos, 2005). These minded rules are used to infer new facts, make logical deductions and answer complex queries. Recently, researchers start to explore combining logical rules into KG embedding to further improve the generalization and performance of HKG reasoning (Alshahrani et al., 2017; Zhu et al., 2022a).
3. Resources
In this section, we compile a detailed resource overview of existing healthcare knowledge graphs to assist researchers and healthcare professionals in constructing and utilizing HKGs, organized in Table 1. We present key attribute information, including HKG name, node types, edge types, statistics, and their applications. Details and external links can be referred to at our public repository [^8] Resource: https://github.com/lujiaying/Awesome-HealthCare-KnowledgeBase.
@@ 3. Resources
Figure 4. Detailed taxonomy of the application section of healthcare knowledge graphs.
4. Applications
4.1. Basic Science Research
Several previous biological terms can also be considered knowledge graphs, such as ontology (gene ontology, cell ontology, disease ontology), network (gene regulatory network), etc. We use the original biological terms as they are more popular according to historical reasons.
4.1.1. Medicinal Chemistry
Topics related to medicinal chemistry involve drug-drug interactions (DDIs) and drug-target interactions (DTIs), which will be discussed in this section.
@@ 4.1.1. Medicinal Chemistry
Drug-drug interactions (DDIs) refer to changes in the actions, or side effects, of drugs when they are taken at the same time or successively (Giacomini et al., 2007). In general, DDIs are a significant contributor to life-threatening adverse events (Su et al., 2022), and their identification is one of the key tasks in public health and drug development. The existence of diverse datasets on drug-drug interactions (DDIs) and biomedical KGs has enabled the development of machine-learning models that can accurately predict DDIs. Yu et al. (Yu et al., 2021) develop SumGNN, a model that includes a subgraph extraction module to efficiently extract relevant subgraphs from a KG, a self-attention-based summarization scheme to generate reasoning paths within the subgraph, and a multichannel module for integrating knowledge and data, resulting in significantly improved predictions of multi-typed DDIs. Su et al. (Su et al., 2022) propose DDKG, an attention-based KG representation learning framework that involves an encoder-decoder layer to learn the initial embeddings of drug nodes from their attributes in the KG. Karim et al. (Karim et al., 2019) compare various techniques for generating KG embeddings with different settings and conclude that a combined convolutional neural network and LSTM network yields the highest accuracy when predicting drug-drug interactions (DDIs). Dai et al. (Dai et al., 2021) propose a new KG embedding framework by introducing adversarial autoencoders based on Wasserstein distances and Gumbel-Softmax relaxation for DDI tasks. Lin et al. (Lin et al., 2020) develop KGNN that resolves the DDI prediction by capturing drugs and their potential neighborhoods by mining associated relations in KG.
@@
Drug-target interactions (DTIs) is just as important as DDIs (Chen et al., 2016). Machine learning models can leverage knowledge graphs constructed from various types of interactions, such as drug-drug, drug-disease, protein-disease, and protein-protein interactions, to aid in predicting DTIs. For instance, Li et al. (Li et al., 2023) utilize the KG transfer probability matrix to redefine the drug-drug and target-target similarity matrix, thus constructing the final graph adjacent matrix to learn node representations by VGAE and augmenting them by utilizing dual Wasserstein Generative Adversarial Network with gradient penalty. Zhang et al. (Zhang et al., 2021c) propose a new hybrid method for DTI prediction by first constructing DTI-related KGs and then employing graph representation learning model to obtain feature vectors of the KG. Wang et al. (Wang et al., 2022a) construct a knowledge graph of 29,607 positive drug-target pairs by DistMult embedding strategy, and propose a Conv-Conv module to extract features of drug-target pairs. Ye et al. (Ye et al., 2021b) learn a low-dimensional representation for various entities in the KG, and then integrate the multimodal information via neural factorization machine.
4.1.2. Bioinformatics Research
In bioinformatics settings, a knowledge graph is a graph-based representation where nodes are biomedical entities (such as mutations, genes, proteins, metabolites, diseases, and biological pathways), and edges are their relationships (such as associations, interactions, and regulations). Through the integration, researchers can gain a more comprehensive understanding of complex biological processes and diseases.
@@ 4.1.2. Bioinformatics Research
Multi-Omics Applications: In recent years, the field of multi-omics data analysis has become increasingly important for understanding complex biological systems. With the advancement of high-throughput technologies, researchers can generate large-scale and high-dimensional data from different omics fields, such as genomics, transcriptomics, proteomics, metabolomics, and epigenomics. More KG applications based on multi-omics data integration have emerged, aiming to provide new research methods to uncover the complex relationships between different omics layers and reveal biological systems’ underlying mechanisms.
@@
Knowledge graphs have been used to identify disease-associated mutations, genes, proteins, and metabolites by integrating multi-omics data with existing biological knowledge. This approach has led to the discovery of novel biomarkers and therapeutic targets for various diseases and the interpretation of the functional effects of genetic elements (Zhang et al., 2021a). Quan et al. built a comprehensive multi-relational knowledge graph called AIMedGraph, providing an interpretation of the impact of genetic variants on disease or treatment (Quan et al., [n. d.]). They curated detailed information about diseases, drugs, genetic variants, and the impact of genetic variations on disease development and drug treatment from multiple data resources. The entities integrated into AIMedGraph are connected by evidence-based relations and form a comprehensive gene–variant–disease–drug–trial–reference knowledge network. AIMedGraph uses the Adamic-Adar algorithm to predict new relations between entities based on shared neighbors and their proximity in the knowledge network. GenomicsKG is a knowledge graph to analyze and visualize multi-omics data. GenomicsKG can be used to improve drug development based on clinical genomics correlations and personalized drug customization in the extended version based on interactive relationships. It also provides multi-dimensional visualization, linked functional knowledge graphs, and reporting for clinical genomics.
@@
Single-Cell Analysis: Cells are fundamental and essential units of living organisms. With high-throughput sequencing technologies advancing to measure genomic profiles in a single-cell resolution, cell functions (inside cells) and cell-cell interactions (between cells) are revealed (Linnarsson and Teichmann, 2016). When diving into the functions of cells, gene regulatory mechanisms can be a critical factor. Gene regulatory mechanisms control the expression of genes, which can affect cell differentiation, response to stimuli, disease progression, etc. It can be visualized as gene regulatory networks (GRNs) which depict the interactions between genes and their regulators. Traditional approaches to constructing the GRN are based on gene knockdown or knockout experiments, which can be time-consuming, labor-intensive, and costly. In addition, these experiments are limited to only one or a few genes which neglects gene-gene interactions. In contrast, single-cell sequencing data could provide whole-genome scale measurements in each individual cell with comprehensive information such as gene expression, transcription factors (TF) binding sites, DNA methylation, epigenetic modifications, etc. By mining these publicly available data, it is possible to reveal the underlying and universal GRN biomedical researchers to understand biological processes better. For example, GRNdb provides detailed regulon and TF-target pairs information from different human and mouse tissues under different conditions by analyzing existing sequencing data (Fang et al., 2021). GenomicKB integrates existing datasets and genome annotations and formulates the data into a knowledge graph to emphasize the relationships among genomic entries (Feng et al., 2023). We can foresee that with more data being generated and collected, one generalized GRN (gene regulatory knowledge graph) along with cell-type-specific GRNs can be reconstructed and used to help reveal the underlying mechanisms of gene expression, biological processes, and disease progression.
Besides cell-type-specific GRN, single-cell sequencing data can also be used to infer cell-cell interactions, which play a critical role in understanding cell cycles, cell fate decisions, tissue development, immune response, etc. Among several approaches through which cells can interact, cell-cell communication or cell signaling is of the most interest (Eltzschig et al., 2006) as the physical distance between cells does not limit it. Cells can communicate with each other by sending signaling molecules called ligands and receiving through receptors located on cell surfaces. Several methods have been developed to implicitly measure cell-cell communication by modeling and scoring ligand-receptor interactions (LRIs) from single-cell sequencing data (Efremova et al., 2020; Jin et al., 2021). However, these methods do not utilize the spatial information of cells. Recent advancements in spatial sequencing techniques provide not only genomic measurements but also complementary information on the spatial coordinates of each cell, where the spatial coordinates indicate the probabilities of cells communicating with each other as cells with closer distances tend to interact more. A recent benchmark study on evaluating cell-cell communication inference methods based on spatial distances (Liu et al., 2022a) while another benchmark study also considers other biological factors such as cytokine activities and receptor protein abundance (Dimitrov et al., 2022). With the availability of spatial transcriptomics data, many methods have been developed to leverage colocalization information to infer LRIs. Although many LRIs databases (Efremova et al., 2020; Shao et al., 2021) have been constructed and applied to infer cell-cell communication, only recently, SpaTalk (Shao et al., 2022) integrates CellTalkDB, Kyoto Encyclopedia of Genes and Genomes (KEGG) pathways, Reactome and TFs from AnimalTFDB to construct a ligand-receptor-target knowledge graph to help improve the inference of cell-cell communication. The success of SpaTalk provides an unprecedented view of integrating existing networks into a knowledge graph and shows that correctly utilizing the knowledge graph could provide helpful information in cell-cell communication tasks.
4.2. Pharmaceutical Research Development
4.2.1. Drug Development
@@
Drug development is the process of identifying novel chemical compounds that can effectively treat or alleviate human diseases. Before the drug can be designated as a final product for clinical use, several critical steps need to be undertaken from the initial target identification, chemical synthesis and clinical trials. The whole process typically spans over a decade and involves expenses of approximately one billion dollars (Adams and Brantner, 2006), yet it is characterized by a low success rate for clinical approval (DiMasi et al., 2015).
With advancements in chemistry and the availability of vast chemical libraries, it is possible to reduce the costs of drug development by extracting and integrating valuable insights from existing data and using computer algorithms to accelerate the drug design process (Du et al., 2021; Zeng et al., 2022; Wang et al., 2022c; Pan et al., 2022b; Wang et al., 2022b; Nicolaou et al., 2007; Nicolaou and Pattichis, 2006). Despite the growing trend of computer-assisted drug discovery, a key question remains regarding how to effectively integrate data and extract valuable insights from the vast chemical dataset.
To approach this question, knowledge graphs (KGs) have been employed for drug discovery due to their various advantages (Zeng et al., 2022): (1) In contrast to traditional methods that capture only one type of relationship, KGs are capable of providing heterogeneous information that includes diverse entities (e.g., scaffolds, proteins, and genes); (2) In addition, KGs are capable of handling multiple types of relationships between various types of entities, such as drug-target pairs; and (3) KGs can provide unstructured semantic relationships between entities. In such graphs, entities are represented as nodes while their relationships are represented as edges, by which complex relations in biochemical systems can be easily handled.
In general, the field of drug development encompasses two main areas: drug design and drug repurposing (also known as drug repositioning). Drug design is the process of creating novel and diverse drug molecules with desirable pharmaceutical properties (Jing et al., 2018; Fu et al., 2021; Fu and Sun, 2022), whereas drug repurposing identifies new uses for existing approved drugs that were originally developed for a different indication (Pan et al., 2022a; Huang et al., 2020).
@@
Drug Design. KGs are widely employed in drug design, particularly in generating novel molecules that are promising drug candidates for various diseases (Ranjan et al., 2022; Li and Chen, 2022). Ranjan et al. (Ranjan et al., 2022) utilize Gated Graph Neural Network (GGNN) to generate novel molecules that target the coronavirus (i.e., SARS-CoV-2) (Hasöksüz et al., 2020) and integrate KGs into their approach to reduce the search space. Specifically, KGs were leveraged to discard non-binding molecules before inputting them into the Early Fusion model, thus optimizing the efficiency of the drug design process. In addition to employing deep learning for direct structure design, KGs are also utilized in the analysis of chemical synthesis. Quantitative estimation of molecular synthetic accessibility is critical in prioritizing the molecules generated from generative models. For instance, Li et al. (Li and Chen, 2022) utilize reaction KGs to construct classification models for compound synthetic accessibility. By leveraging KGs that capture information about reactions, including reaction types, substrates, and reaction conditions, they can train machine learning models that could predict the synthetic accessibility of compounds. Jeong et al. (Jeong et al., 2022) introduce an intelligent system that integrates generative exploration and exploitation of reaction knowledge base to support synthetic path design.
@@
Drug Repurposing. Compared to drug design, KGs are more commonly utilized to expedite the drug re-purposing process (Zhu et al., 2020; MacLean, 2021; Himmelstein et al., 2017; Zhang et al., 2021b; Gao et al., 2022; Xu et al., 2019; Zhang and Che, 2021; Ghorbanali et al., 2023). Many applications on drug re-purposing that utilize KGs are primarily focused on link prediction tasks (MacLean, 2021). To re-purpose promising drug candidates for new indications, many methods employ predictive models that focus on predicting drug-treats-disease relationships within pharmacological knowledge graphs KGs. Himmelstein et al. (Himmelstein et al., 2017) use a degree-normalized pathway model on the hetionet KG, which includes genes, diseases, tissues, pathophysiologies, and multimodal edges, to identify potentially repurposable drugs for epilepsy. Xu et al. (Xu et al., 2019) develop a multi-path random walk model on a network that incorporates gene-phenotype associations, protein-protein interactions, and phenotypic similarities for training and prediction purposes. Zhang et al. (Zhang et al., 2021b) introduce an integrative and literature-based discovery model for identifying potential drug candidates from COVID-19-focused research literature, including PubMed and other relevant sources. Gao et al. (Gao et al., 2022) construct a knowledge graph (KG) by integrating multiple genotypic and phenotypic databases. They then learn low-dimensional representations of the KG and utilize these representations to infer new drug-disease interactions, providing insights into potential drug repurposing opportunities. Zhang and Che (Zhang and Che, 2021) introduce a model for drug re-purposing in Parkinson’s disease that leverages a local medical knowledge base incorporating accurate knowledge along with medical literature containing novel information. Ghorbanali et al. (Ghorbanali et al., 2023) present the DrugRep-KG method, which utilizes a KG embedding approach for representing drugs and diseases in the process of drug repurposing.
4.2.2. Clinical Trial
@@
The major goal of clinical trials is to assess the safety and effectiveness of drug molecules on human bodies. A novel drug molecule needs to pass three phases of clinical trials before it is approved by the Food and Drug Administration (FDA) and enters the drug market. The whole process is prohibitively time-consuming and expensive, costing 7-11 years and two billion dollars on average (Martin et al., 2017).
@@
Clinical Trial Optimization targets identifying eligible patients for clinical trials based on their medical history and health conditions (Rivera et al., 2020; He et al., 2020). Recently, with massive electronic health records (EHR) data and trial eligibility criteria (EC), data-driven methods have been studied to automatically assign appropriate patients for clinical trials (Yuan et al., 2019; Tseo et al., 2020; Liu et al., 2021). However, it is often hard to fully capture and represent the complex knowledge present in unstructured ECs and EHR data, as ECs may only provide general disease concepts. In contrast, patient EHR data contain more specific medical codes to represent patient conditions. To better capture the interactions among different medical concepts from EHR records and ECs, Gao et al. (Gao et al., 2020) enhance patient records with hierarchical taxonomies to align medical concepts of varying granularity between EHR codes and ECs. Besides, Fu et al. (Fu et al., 2022) leverage additional knowledge-embedding modules along with drug pharmacokinetic and historical trial data to improve the patient trial optimization process, and Wang et al. (Wang et al., 2023) leverage the knowledge graphs to learn static trial embedding and further designed meta-learning module to generalize well over the imbalanced clinical trial distribution.
4.3. Clinical Decision Support
Nowadays, abundant Electronic Health Record (EHR) data enables better computational models for accurate diagnoses and treatments. EHR contains essential patient information such as disease diagnoses, prescribed medications, and test results. Due to this valuable information, EHRs are extensively utilized to identify patterns in patient health and assist healthcare providers in making informed clinical decisions.

### v5

@@ 2.2 HKG Utilization
Model-free Utilization. Various query languages can be used for KGs, such as SPARQL, Cypher, and GraphQL 44. These query languages allow users to query HKGs using a standardized syntax, thus enabling users to retrieve, manipulate, and analyze data in a structured and consistent way. For instance, automatic healthcare question answering can be tackled by Natural Language Question-to-Query (NLQ2Query) approach 97, where natural language questions are first translated into executable graph queries and then answered by the query responses. HKGs can also be utilized as an up-to-date and trustworthy augmentation to LLMs for many applications. Some pioneering studies 98; 99; 100 show that retrieved knowledge triples can improve the reliability of LLMs in various knowledge-intensive tasks. Moreover, KGs can be a useful tool for fact-checking 101; 102; 103 as they provide a structured representation of information that can be used to quickly and efficiently verify the accuracy of claims. Researchers have explored the utility of HKGs in identifying ingredient substitutions of food 104, COVID-19 fact-checking 105, etc.
@@
Model-based Utilization. Utilizing HKGs in complex reasoning tasks often involves utilizing machine learning models. HKG embeddings 106; 107 have shown great potential to tackle these tasks. In particular, HKG embedding models are a class of machine learning models that aim to learn low-dimensional vector representations of the entities and relations in a KG. After obtaining HGK embeddings, they can be plugged into any kind of deep neural network and further fine-tuned toward downstream objectives. On the other hand, Symbolic logic models offer an interpretable approach to KG reasoning by mining logical rules from existing knowledge through techniques such as inductive logic programming 108, association rule mining 109, or Markov logic networks 110. These minded rules are used to infer new facts, make logical deductions and answer complex queries. Recently, researchers start to explore combining logical rules into KG embedding to further improve the generalization and performance of HKG reasoning 111; 112.
3 Applications
@@
Figure 5: Detailed taxonomy of the application section of healthcare knowledge graphs.
3.1 Basic Science Research
Biologists have built various ontologies (gene, cell, disease) and networks (gene regulatory, protein interaction) over the past decade, which are also KGs. They are commonly used in medicinal chemistry and bioinformatic research.
3.1.1 Medicinal Chemistry
@@
Drug-drug interactions (DDIs) refer to changes in the actions, or side effects, of drugs when they are taken at the same time or successively 113. In general, DDIs are a significant contributor to life-threatening adverse events 107, and their identification is one of the key tasks in public health and drug development. The existence of diverse datasets on drug-drug interactions (DDIs) and biomedical KGs has enabled the development of machine-learning models that can accurately predict DDIs. Yu et al. 106 develop SumGNN, a model that efficiently extract relevant subgraphs from a KG to generate reasoning paths within the subgraph, resulting in significantly improved predictions of multi-typed DDIs. Su et al. 107 propose DDKG, an attention-based KG representation learning framework that involves an encoder-decoder layer to learn the initial embeddings of drug nodes from their attributes in the KG. Karim et al. 114 compare various techniques for generating KG embeddings with different settings and conclude that a combined convolutional neural network and LSTM network yields the highest accuracy when predicting drug-drug interactions (DDIs). Dai et al. 115 propose a new KG embedding framework by introducing adversarial autoencoders based on Wasserstein distances and Gumbel-Softmax relaxation for DDI tasks. Lin et al. 116 develop KGNN that resolves the DDI prediction by capturing drugs and their potential neighborhoods by mining associated relations in KG.
@@
Drug-target interactions (DTIs) are as important as DDIs 117. Machine learning models can leverage KGs constructed from various types of interactions, such as drug-drug, drug-disease, protein-disease, and protein-protein interactions, to aid in predicting DTIs. For instance, Li et al. 118 utilize the KG transfer probability matrix to redefine the drug-drug and target-target similarity matrix, thus constructing the final graph adjacent matrix to learn node representations by utilizing dual Wasserstein Generative Adversarial Network. Zhang et al. 119 propose a new hybrid method for DTI prediction by first constructing DTI-related KGs and then employing graph representation learning model to obtain feature vectors of the KG. Wang et al. 120 construct a KG of 29,607 positive drug-target pairs by DistMult embedding strategy, and propose a Conv-Conv module to extract features of drug-target pairs. Ye et al. 121 learn a low-dimensional representation for various entities in the KG, and then integrate the multimodal information via neural factorization machine.
3.1.2 Bioinformatics Research
@@
Multi-Omics Analysis has become increasingly important for understanding complex biological systems. With the advancement of high-throughput technologies, more KG applications based on multi-omics data integration have emerged, aiming to provide new research methods to uncover the complex relationships between different omics layers and reveal biological systems’ underlying mechanisms.
@@
KGs have been used to identify disease-associated mutations, genes, proteins, and metabolites by integrating multi-omics data with existing biological knowledge. This approach has led to the discovery of novel biomarkers and therapeutic targets for various diseases 122. Quan et al. built a comprehensive multi-relational KG called AIMedGraph, providing an interpretation of the impact of genetic variants on disease or treatment 123. They curated detailed information about diseases, drugs, genetic variants, and the impact of genetic variations on disease development and drug treatment from multiple data resources. GenomicsKG is a KG to analyze and visualize multi-omics data. GenomicsKG can be used to improve drug development based on clinical genomics correlations and personalized drug customization in the extended version based on interactive relationships. It also provides multi-dimensional visualization, linked functional KGs, and reporting for clinical genomics.
@@
Single-Cell Analysis focuses on cells as the fundamental and essential units of living organisms. With high-throughput sequencing technologies advancing to measure genomic profiles in a single-cell resolution, cell functions (inside cells) and cell-cell interactions (between cells) are revealed 124. Gene regulatory mechanisms, which control gene expression and affect processes like cell differentiation and disease progression, are key to understanding these functions. Traditional gene knockdown experiments are time-consuming and limited, but single-cell sequencing provides genome-wide data, including gene expression, transcription factor binding, DNA methylation, and epigenetic modifications. This allows researchers to uncover gene regulatory networks (GRNs) that enhance our understanding of biological processes. Databases like GRNdb and GenomicKB integrate sequencing data and annotations to provide insights into gene regulation across tissues 125; 126, which will improve as more data becomes available.
3.2 Pharmaceutical Research Development
3.2.1 Drug Development
@@
Drug development is the process of identifying novel chemical compounds that can effectively treat or alleviate human diseases. Before the drug can be designated as a final product for clinical use, several critical steps need to be undertaken from the initial target identification, chemical synthesis and clinical trials. The whole process typically spans over a decade and involves expenses of approximately one billion dollars 127, yet it is characterized by a low success rate for clinical approval 128.
@@
Drug Design is an area where KGs are commonly used, particularly for generating novel molecules that are promising drug candidates for various diseases 86; 87. Ranjan et al. 86 utilize Gated Graph Neural Network (GGNN) to generate novel molecules that target the coronavirus (i.e., SARS-CoV-2) 129 and integrate KGs into their approach to reduce the search space. Specifically, KGs were leveraged to discard non-binding molecules before inputting them into the Early Fusion model, thus optimizing the efficiency of the drug design process. In addition to employing deep learning for direct structure design, KGs are also utilized in the analysis of chemical synthesis. Quantitative estimation of molecular synthetic accessibility is critical in prioritizing the molecules generated from generative models. For instance, Li et al. 87 utilize reaction KGs to construct classification models for compound synthetic accessibility. By leveraging KGs that capture information about reactions, including reaction types, substrates, and reaction conditions, they can train machine learning models that could predict the synthetic accessibility of compounds. Jeong et al. 88 introduce an intelligent system that integrates generative exploration and exploitation of reaction knowledge base to support synthetic path design.
@@
Drug Repurposing has often been expedited by the utilization of KGs 130; 131; 47; 89; 90; 132; 91; 92. Many applications on drug re-purposing that utilize KGs are primarily focused on link prediction tasks 131. To re-purpose promising drug candidates for new indications, many methods employ predictive models that focus on predicting drug-treats-disease relationships within pharmacological KGs. Xu et al. 132 develop a multi-path random walk model on a network that incorporates gene-phenotype associations, protein-protein interactions, and phenotypic similarities for training and prediction purposes. Zhang et al. 89 introduce an integrative and literature-based discovery model for identifying potential drug candidates from COVID-19-focused research literature, including PubMed and other relevant sources. Gao et al. 90 construct a KG by integrating multiple genotypic and phenotypic databases. They then learn low-dimensional representations of the KG and utilize these representations to infer new drug-disease interactions, providing insights into potential drug repurposing opportunities. Zhang and Che 91 introduce a model for drug re-purposing in Parkinson’s disease that leverages a local medical knowledge base incorporating accurate knowledge along with medical literature containing novel information. Ghorbanali et al. 92 present the DrugRep-KG method, which utilizes a KG embedding approach for representing drugs and diseases in the process of drug repurposing.
3.2.2 Clinical Trial
@@
The major goal of clinical trials is to assess the safety and effectiveness of drug molecules on human bodies. A novel drug molecule needs to pass three phases of clinical trials before it is approved by the Food and Drug Administration (FDA) and enters the drug market. The whole process is prohibitively time-consuming and expensive, costing 7-11 years and two billion dollars on average 133.
@@
Clinical Trial Optimization targets identifying eligible patients for clinical trials based on their medical history and health conditions 134; 135. Recently, with massive electronic health records (EHR) data and trial eligibility criteria (EC), data-driven methods have been studied to automatically assign appropriate patients for clinical trials 136; 137; 138. However, it is often hard to fully capture and represent the complex knowledge present in unstructured ECs and EHR data, as ECs may only provide general disease concepts. In contrast, patient EHR data contain more specific medical codes to represent patient conditions. To better capture the interactions among different medical concepts from EHR records and ECs, Gao et al. 139 enhance patient records with hierarchical taxonomies to align medical concepts of varying granularity between EHR codes and ECs. Besides, Fu et al. 140 leverage additional knowledge-embedding modules along with drug pharmacokinetic and historical trial data to improve the patient trial optimization process, and Wang et al. 141 leverage the KGs to learn static trial embedding and further designed meta-learning module to generalize well over the imbalanced clinical trial distribution.
3.3 Clinical Decision Support